KjSaif:

The Soccer Club is selling T-shirts and washing cars to raise money for a summer trip. Last week, the club raised $292.50 by selling 15 T-shirts and washing 12 cars. This week, it raised $747.50 by selling 45 T-shirts and washing 10 cars. Write two equations showing the money the club raised each week, where x is the amount earned per T-shirt sold and y is the amount earned per car washed. How much does the Soccer Club charge for a T-shirt? Show your work. How much does the Soccer Club charge for a car wash? Show your work.

KjSaif:

@WillieXO

Shadow:

Let t = amount charged for t shirt Let c = amount charged for cars 15t + 12c = 292.50 45t + 10c = 747.50

KjSaif:

thanks! so...know it's just saying ............ How much does the Soccer Club charge for a T-shirt? Show your work. How much does the Soccer Club charge for a car wash? Show your work.

KjSaif:

@shadow

WillieXO:

15x + 12y = 292.50 45x + 10y = 747.50 use subtiutution method or elimination whatever is best i would use elimination and multiply the 1st equation by -3 therefor -3(15x + 12y) = 292.50 45x + 10y = 747.50 simplify and solve

KjSaif:

hmmmm....im kinda stuck....give me a sec @WillieXO

KjSaif:

does the car wash cost $3.61?

KjSaif:

and do the t shirts cost $16.61?

KjSaif:

sorry if i'm wrong ;-;

WillieXO:

-3(15x + 12y) = 292.50 45x + 10y = 747.50 ================= -45x - 36y = -877.5 45x + 10y = 747.50 ============= -26y = -130 y = 5 now let's find x 45x + 10(5) = 747.50 45x + 50 = 747.50 x = 15.5 So the T shirts 15.5 and cars 5

KjSaif:

thanks! :)

WillieXO:

welcome

KjSaif:

i knew i was off somewhere xD

WillieXO:

xD

Shadow:

Sorry, had to go. Glad willie was here

WillieXO:

Thanks shadow